OpenStudy (anonymous):

3x+1=2x+2 SOLVE FOR X!!!!

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Can you help explain please?

OpenStudy (anonymous):

3x+1=2x+2 -1 -1 3x=2x+1 -2x -2x x = 1 3(1)+1=4 2(1) +2=4 4=4

OpenStudy (anonymous):

Thank youuuu!!!!

OpenStudy (anonymous):

Can you answer another please?

OpenStudy (anonymous):

Write an indirect proof to show that a rectangle has congruent diagonals. Be sure to create and name the appropriate geometric figures. This figure does not need to be submitted.